A person pulls a toboggan for a distance of 30.3 m along the snow with a rope directed 32.8 ° above the snow. The tension in the rope is 102 N. (a) How much work is done on the toboggan by the tension force? (b) How much work is done if the same tension is directed parallel to the snow?
Answer #1

here,

the distance travelled , s = 30.3 m

tension in the rope , T = 102 N

a)

theta = 32.8 degree

the work done on the toboggan by the tension force , W = T * s * cos(theta)

W = 102 * 30.3 * cos(32.8) J

W = 2597.86 J

b)

theta = 0 degree

the work done on the toboggan by the tension force , W = T * s * cos(0)

W = 102 * 30.3 * cos(0) J

W = 3090.6 J